# Geometry of quantum states and chaos-integrability transition

### v1

@@ Acknowledgments
It is a pleasure to thank Hugo A. Camargo, Yichao Fu, and Viktor Jahnke for many insightful discussions and for comments on a draft version of the manuscript. The work of Kunal Pal is supported by the YST Program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government. This was also supported by the Korean Local Governments - Gyeongsangbuk-do Province and Pohang City. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791, NRF-2020R1I1A2054376), the framework of international cooperation program managed by the NRF of Korea (RS-2023-NR119907, RS-2024-00440725), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the NRF funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025.
@@ Appendix A QGT for a generic superposition of eigenstates
$g_{ab}(\mathbf{r})=\braket{\partial_{a}\Psi(\mathbf{r})}{\partial_{b}\Psi(\mathbf{r})}-\braket{\partial_{a}\Psi(\mathbf{r})}{\Psi(\mathbf{r})}\braket{\Psi(\mathbf{r})}{\partial_{b}\Psi(\mathbf{r})}~{},$ (83)
for a generic superposition of eigenstates, |$\ket{\Psi(\mathbf{r})}=\sum_{n}c_{n}\ket{n(\mathbf{r})}$, where $c_{n}$s are complex numbers independent of the set of parameters $\{\mathbf{r}\}$ present in the total Hamiltonian. A straightforward calculation shows that the expression for the QGT can be written as
$\begin{split}g_{ab}(\mathbf{r})=\sum_{n,m}c_{n}^{*}c_{m}~{}\Bigg{(}\sum_{k}(1-|c_{k}|^{2})\braket{\partial_{a}n(\mathbf{r})}{k(\mathbf{r})}\braket{k(\mathbf{r})}{\partial_{b}m(\mathbf{r})}\\ -\sum_{k\neq l}c_{k}c_{l}^{*}\braket{\partial_{a}n(\mathbf{r})}{k(\mathbf{r})}\braket{l(\mathbf{r})}{\partial_{b}m(\mathbf{r})}\Bigg{)}~{}.\end{split}$ (84)
Specifically, for the state |$\ket{\Psi(\mathbf{r})}=N^{-1/2}\sum_{n}\ket{n(\mathbf{r})}$, the components of QGT are given by
$\begin{split}g_{ab}(\mathbf{r})=\frac{1}{N}\sum_{n,m}~{}\Bigg{(}\sum_{k}(1-\frac{1}{N})\braket{\partial_{a}n(\mathbf{r})}{k(\mathbf{r})}\braket{k(\mathbf{r})}{\partial_{b}m(\mathbf{r})}\\ -\sum_{k\neq l}\frac{1}{N}\braket{\partial_{a}n(\mathbf{r})}{k(\mathbf{r})}\braket{l(\mathbf{r})}{\partial_{b}m(\mathbf{r})}\Bigg{)}~{}.\end{split}$ (85)
@@ Appendix B Derivation of QMT component for GUE Hamiltonian
In this Appendix, we briefly present the derivation of the $rr$ component of the ensemble-averaged QMT ($\bar{g}_{rr}$ in (24)) for the GUE Hamiltonian in (21).181818 We note that the derivation presented in [42] contains a few typographical errors. We start by writing the first expression in (22) explicitly (since $H=H_{0}+r\tilde{\mathcal{H}}_{1}$),
@@
where we have performed the trivial integration over ~$\tilde{\mathcal{H}}_{2}$. Now performing two successive variable changes: first using $H_{0}=H-r\tilde{\mathcal{H}}_{1}$ and subsequently ~$\tilde{\mathcal{H}}_{1}=\frac{H_{1}}{\sqrt{r^{2}+1}}+\frac{r}{r^{2}+1}H$, we can rewrite the above expression as, [^19] All the Jacobian factors associated with these transformations are constant, and hence cancel from the same factors form the integral in $Z$ in the denominator.
@@
Note that the motivation behind these particular coordinate changes is to write the total probability distribution in a form which factorises in two variables, one of which should be the total Hamiltonian $H$. Also note that, with the conventions fixed at the beginning of section III, the variance of the variable $H_{1}$ is $\sigma^{2}$. We can perform the integral over $H_{1}$ by using the standard formula for the correlator of the GUE matrices, $\mathbb{E}\big{(}(H_{1})_{ij}(H_{1})_{kl}\big{)}=\sigma^{2}\delta_{il}\delta_{jk}$. Since we have fixed $\sigma^{2}=1/N$, we then have,
@@ Appendix C Coordinate transformations to the metric of $S^{2}$
Here, we briefly discuss the coordinate transformation, which maps the line element into the round metric on the 2-sphere $S^{2}$. Taking $r^{2}=\cot^{2}\theta$, with $\theta\in[-\pi/2,\pi/2]$, we get the transformed line element (with $A=(N-1)/2$)
@@ Appendix D Alternative expression for the correlation function with GUE matrix
In this section, we discuss an alternative expression for the correlation function in eq. (36) that will illustrate some important points regarding how these correlation functions evolve with time and the role of the rank of the Hamiltonian $N$, specifically the limit $N\rightarrow\infty$.
To compute the required correlation function, in the following, we follow an approach inspired by the free probability theory [48, 73, 47]. The basic idea is the following. Since the random matrices $H_{0}$ and $V$ are independent, in the limit of large $N\rightarrow\infty$, they become free random variables. However, since in the correlation function the time evolution is generated by the total Hamiltonian $\tilde{H}$, what we need is a random variable which is free from $\tilde{H}$. One way of doing this is to find out an orthogonal transformation which transforms the free variables $H_{0}$ and $V$ to the new free variables $\tilde{H}$ and, say, $X$. After obtaining $X$ in terms of $H_{0}$ and $V$, one can transform moments of the form $\braket{\tilde{H}V(t)\tilde{H}V}$ into moments in terms of mutually free variables, as, $\braket{\tilde{H}X(t)\tilde{H}X}$, which can then be computed using free probability. [^20] In the following we follow an approach outlined in [74].
@@
where $\alpha=\sqrt{(r^{2}+1)}\sigma$. Note that the expression for the random matrix $X$ is the same as the one obtained in (39) by requiring that the total probability distribution factorises in terms of the variables $\tilde{H}$ and $X$ another way of getting independent Gaussian variables. Since $X$ and $H$ are independent variables, drawn from Gaussian ensembles, in the limit $N\rightarrow\infty$, they would be free variables. Now we can rewrite the correlator $\braket{V(t)V}$ under consideration as follows
@@
Using the fact that $\braket{X}=\braket{\tilde{H}}=0$, we see that $\braket{X\tilde{H}}=0$ from the freeness of these two variables. Furthermore, the correlator $\braket{X(t)X}$ can be expanded in a Taylor series with terms of the form $\braket{\tilde{H}^{i}X\tilde{H}^{j}X}$. Once again, using the property of free variables that the crossing partitions do not contribute in the moment expansion, these moments can be evaluated by noting that the only diagram that has a non-zero contribution gives a result of the form $\braket{\tilde{H}^{i}}\braket{X^{2}}\braket{\tilde{H}^{j}}$. Using all these inputs in eq. (95) we get the final form of the desired correlator to be (using $\braket{X^{2}}=N$)
@@
Comparing this expression for the correlator (in (97)) with the earlier expression derived in (42), we see that, in this expression, the time dependence comes from only the connected part of the SFF, whereas, in (42), the full SFF contributes to the correlation function. This is essentially related to the fact that in the large $N$ limit all the terms in the SFF, apart from the one coming the disconnected part of the two-point density correlation fucntion vanish in the $N\rightarrow\infty$ limit.
@@
Here $|\braket{e^{-it^{\prime}H^{\prime}}}|^{2}$ denotes the contribution of the disconnected piece of the $2$-point correlation function, while $r_{2}(t^{\prime})$ is the contribution of the connected part that gives rise to the ramp in the SFF. Now, due to the $N^{2}$ factor present in the denominator of (43), we see from the above expression for $\mathcal{R}(t^{\prime})$ that both the constant term and the contribution coming from the connected part ($r_{2}(t^{\prime})$) are suppressed as $1/N$ in the limit $N\rightarrow\infty$, and the only contribution in this limit comes from $|\braket{e^{-it^{\prime}H^{\prime}}}|^{2}$. Note that the function $r_{2}(t^{\prime})$ is only non-zero up to a timescale $t^{\prime}<2N$, hence even though this term is linearly increasing, in the late times $t^{\prime}>2N$ it does not contribute to the correlation function. In the free probabilistic computation, in eq. (95), this can be traced back to the fact that only non-crossing partitions contribute to the moments [75].
@@ Appendix E Computation of the fidelity susceptibility for a $2\times 2$ random matrix ensemble
In this Appendix, we present a random matrix ensemble with generic Dyson index $\beta$, which we use to compute the FS, both for the case of ensembles with only chaotic phase, and also one showing a chaos-to-integrability transition.
@@
In the following, we assume that the joint eigenvalues of the matrices $H_{0}$ and $V$ are independent, and are drawn from the joint probability distribution (6) of the Gaussian $\beta$-ensemble. The total Hamiltonian of the system we consider is given by
@@
Case 1. First, consider the case where the joint probability distributions of both $h_{i}$ and $v_{i}$ are the same, i.e., the ensembles from which these matrices are drawn have the same Dyson index. The ensemble-averaged FS computed from this Hamiltonian is given by (with $h_{12}=h_{1}-h_{2}$ and $v_{12}=v_{1}-v_{2}$)

### v2

@@ Acknowledgments
It is a pleasure to thank Hugo A. Camargo, Yichao Fu, and Viktor Jahnke for many insightful discussions and for comments on a draft version of the manuscript. The work of Kunal Pal is supported by the YST Program at the APCTP through the Science and Technology Promotion Fund and Lottery Fund of the Korean Government. This was also supported by the Korean Local Governments - Gyeongsangbuk-do Province and Pohang City. Kunal Pal would like to thank the hospitality of the Gwangju Institute of Science and Technology for the hospitality where a part of this work was carried out. This work was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Science, ICT & Future Planning (NRF-2021R1A2C1006791, NRF-2020R1I1A2054376), the framework of international cooperation program managed by the NRF of Korea (RS-2023-NR119907, RS-2024-00440725), the Creation of the Quantum Information Science R&D Ecosystem (Grant No. 2022M3H3A106307411) through the NRF funded by the Korean government (Ministry of Science and ICT), the Gwangju Institute of Science and Technology (GIST) research fund (Future leading Specialized Resarch Project, 2025) and the Al-based GIST Research Scientist Project grant funded by the GIST in 2025.
@@ Appendix A QGT for a generic superposition of eigenstates
$g_{ab}(\mathbf{r})=\braket{\partial_{a}\Psi(\mathbf{r})|\partial_{b}\Psi(\mathbf{r})}-\braket{\partial_{a}\Psi(\mathbf{r})|\Psi(\mathbf{r})}\braket{\Psi(\mathbf{r})|\partial_{b}\Psi(\mathbf{r})}~,$ (96)
[…]
$\begin{split}g_{ab}(\mathbf{r})=\sum_{n,m}c_{n}^{*}c_{m}~\Bigg{(}\sum_{k}(1-|c_{k}|^{2})\braket{\partial_{a}n(\mathbf{r})|k(\mathbf{r})}\braket{k(\mathbf{r})|\partial_{b}m(\mathbf{r})}\\ -\sum_{k\neq l}c_{k}c_{l}^{*}\braket{\partial_{a}n(\mathbf{r})|k(\mathbf{r})}\braket{l(\mathbf{r})|\partial_{b}m(\mathbf{r})}\Bigg{)}~.\end{split}$ (97)
[…]
$\begin{split}g_{ab}(\mathbf{r})=\frac{1}{N}\sum_{n,m}~\Bigg{(}\sum_{k}(1-\frac{1}{N})\braket{\partial_{a}n(\mathbf{r})|k(\mathbf{r})}\braket{k(\mathbf{r})|\partial_{b}m(\mathbf{r})}\\ -\sum_{k\neq l}\frac{1}{N}\braket{\partial_{a}n(\mathbf{r})|k(\mathbf{r})}\braket{l(\mathbf{r})|\partial_{b}m(\mathbf{r})}\Bigg{)}~.\end{split}$ (98)
@@ Appendix B Derivation of QMT component for GUE Hamiltonian
In this Appendix, we briefly present the derivation of the $rr$ component of the ensemble-averaged QMT ($\bar{g}_{rr}$ in (27)) for the GUE Hamiltonian in (24).212121 We note that the derivation presented in [44] contains a few typographical errors. We start by writing the first expression in (25) explicitly (since $H=H_{0}+r\tilde{\mathcal{H}}_{1}$),
@@
where we have performed the trivial integration over $\tilde{\mathcal{H}}_{2}$. Now performing two successive variable changes: first using $H_{0}=H-r\tilde{\mathcal{H}}_{1}$ and subsequently $\tilde{\mathcal{H}}_{1}=\frac{H_{1}}{\sqrt{r^{2}+1}}+\frac{r}{r^{2}+1}H$, we can rewrite the above expression as, [^22] All the Jacobian factors associated with these transformations are constant, and hence cancel from the same factors form the integral in $Z$ in the denominator.
@@
[…]
@@ Appendix C Coordinate transformations to the metric of $S^{2}$
Here, we briefly discuss the coordinate transformation, which maps the line element into the round metric on the 2-sphere $S^{2}$. Taking $r^{2}=\cot^{2}\theta$, with $\theta\in[-\pi/2,\pi/2]$, we get the transformed line element (with A=(N1)/2A=(N-1)/2)
@@ Appendix D Alternative expression for the correlation function with GUE matrix
In this appendix, we obtain an alternative expression for the correlation function in eq. (39) using techniques of free probability theory that will illustrate some important points regarding how these correlation functions evolve with time and the role of the rank of the Hamiltonian $N$, specifically the limit $N\rightarrow\infty$.
To compute the required correlation function, in the following, we follow an approach inspired by the free probability theory [54, 86, 53]. The basic idea is the following. Since the random matrices $H_{0}$ and $V$ are independent [^23] We are considering the scaled Hamiltonian setting described at the beginning of section IV., in the limit of large $N\rightarrow\infty$, they become free random variables. However, since in the correlation function the time evolution is generated by the total Hamiltonian $\tilde{H}$, what we need is a random variable which is free from $\tilde{H}$. One way of doing this is to find out an orthogonal transformation which transforms the free variables $H_{0}$ and $V$ to the new free variables $\tilde{H}$ and, say, $X$. After obtaining $X$ in terms of $H_{0}$ and $V$, one can transform moments of the form $\braket{\tilde{H}V(t)\tilde{H}V}$ into moments in terms of mutually free variables, as, $\braket{\tilde{H}X(t)\tilde{H}X}$, which can then be computed using free probability. [^24] In the following we follow an approach outlined in [87].
@@
where $\alpha=\sqrt{(r^{2}+1)}\sigma$. Note that the expression for the random matrix $X$ is the same as the one obtained in (42) by requiring that the total probability distribution factorises in terms of the variables $\tilde{H}$ and $X$ - another way of getting independent Gaussian variables. Since $X$ and $\tilde{H}$ are independent variables, drawn from Gaussian ensembles, in the limit $N\rightarrow\infty$, they would be free variables. Now we can rewrite the correlator $\braket{V(t)V}$ under consideration as follows
@@
Using the fact that $\braket{X}=\braket{\tilde{H}}=0$, we see that $\braket{X\tilde{H}}=0$ from the freeness of these two variables. Furthermore, the correlator $\braket{X(t)X}$ can be expanded in a Taylor series in terms of the mixed moments of the form $\braket{\tilde{H}^{i}X\tilde{H}^{j}X}$. Once again, using the property of free variables that the crossing partitions do not contribute in the moment expansion, these moments can be evaluated by noting that the only diagram that has a non-zero contribution gives a result of the form $\braket{\tilde{H}^{i}}\braket{X^{2}}\braket{\tilde{H}^{j}}$. Using all these inputs in eq. (108) we get the final form of the desired correlator to be (using $\braket{X^{2}}=N$)
@@
Comparing this expression for the correlator (in (110)) with the earlier expression derived in (45), we see that, in this expression, the time dependence comes from only the connected part of the SFF, whereas, in (45), the full SFF contributes to the correlation function. This is essentially a consequence of the fact that in the large $N$ limit all the terms in the SFF, apart from the one coming from the disconnected part of the two-point density correlation function, vanish in the $N\rightarrow\infty$ limit.
@@
Here $|\braket{e^{-it^{\prime}H^{\prime}}}|^{2}$ denotes the contribution of the disconnected piece of the $2$-point correlation function, while $r_{2}(t^{\prime})$ is the contribution of the connected part that gives rise to the ramp in the SFF. Now, due to the $N^{2}$ factor present in the denominator of (46), we see from the above expression for $\mathcal{R}(t^{\prime})$ that both the constant term and the contribution coming from the connected part ($r_{2}(t^{\prime})$) are suppressed as $1/N$ in the limit $N\rightarrow\infty$, and the only contribution in this limit comes from $|\braket{e^{-it^{\prime}H^{\prime}}}|^{2}$. Note that the function $r_{2}(t^{\prime})$ is only non-zero up to a timescale $t^{\prime}<2N$, hence even though this term is linearly increasing, in the late times $t^{\prime}>2N$ it does not contribute to the correlation function. In the free probabilistic computation, in eq. (108), this can be traced back to the fact that only non-crossing partitions contribute to the moments (see [88] for a related discussion on a similar observation about four-point out-of-time ordered correlators of spin operators).
@@ Appendix E Computation of the fidelity susceptibility for a $2\times 2$ random matrix ensemble
In this Appendix, we consider a random matrix ensemble, an element of which has the eigenvalue distribution of that of a Gaussian ensemble with generic Dyson index $\beta$ and has Haar random eigenstates, and we use it to compute the FS, both for the case of ensembles with only chaotic phase, and also one showing a chaos-to-integrability transition.
@@
In the following, we assume that the joint distributions of eigenvalues of the matrices $H_{0}$ and $V$ are independent, and are drawn from the joint probability distribution (6) of the Gaussian $\beta$-ensemble. The total Hamiltonian of the system we consider is given by
@@
Case 1. First, consider the case where the JPDs of both $h_{i}$ and $v_{i}$ are the same, i.e., the ensembles from which these matrices are drawn have the same Dyson index. The ensemble-averaged FS computed from this Hamiltonian is given by (with $h_{12}=h_{1}-h_{2}$ and $v_{12}=v_{1}-v_{2}$)